The following table shows price and quantity data of Country BF which produces only two products: food and beverages.

Beverages products

Food products

Price $

Quantity (tons)

Price $

Quantity (tons)

2017

2

20

10

10

2018

3

30

11

20

2019

4

35

13

25

    Question:  the inflation rate for 2018 and 2019 using the GDP deflator
Answer #1

As it is not mentioned in the question, let us assume the base year to be 2017

As 2017 is the base year, GDP deflator in 2017 = 100

Nominal GDP in 2018 = Quantity of beverages in 2018 x Price of beverages in 2018 + Quantity of food products in 2018 x Price of food products in 2018 = 30 x $3 + 20 x $11 = $90 + $220 = $310

Real GDP in 2018 = Quantity of beverages in 2018 x Price of beverages in 2017 + Quantity of food products in 2018 x Price of food products in 2017 = 30 x $2 + 20 x $10 = $60 + $200 = $260

GDP deflator in 2018 = [Nominal GDP in 2018/Real GDP in 2018] x 100 = [$310/$260] x 100 = 119.23

Nominal GDP in 2019 = Quantity of beverages in 2019 x Price of beverages in 2019 + Quantity of food products in 2019 x Price of food products in 2019 = 35 x $4 + 25 x $13 = $140 + $325 = $465

Real GDP in 2019 = Quantity of beverages in 2019 x Price of beverages in 2017 + Quantity of food products in 2019 x Price of food products in 2017 = 35 x $2 + 25 x $10 = $70 + $250 = $320

GDP deflator in 2019 = [Nominal GDP in 2019/Real GDP in 2019] x 100 = [$465/$320] x 100 = 145.31

Rate of inflation in 2018 = [(GDP Deflator in 2018 - GDP Deflator in 2017)/GDP Deflator in 2017] x 100 = [(119.23-100)/100] x 100 = 19.23/100 x 100 = 19.23%

Rate of inflation in 2019 = [(GDP Deflator in 2019 - GDP Deflator in 2018)/GDP Deflator in 2018] x 100 = [(145.31-119.23/119.23] x 100 = [26.08/119.23] x 100 = 21.87%